The six function values for 4/5 + Pi/2
Answer by lwsshak3(11628)

The six function values for 4/5 + Pi/2
4/5+π/2=0.8+π/2=2.371 radians (In quadrant II)
Use calculator set to radians:
sin=0.697
cos=-0.717
tan=-0.971
csc=1/sin=1.435
sec=1/cos=-1.395
cot=1/tan=-1.030